Calculate the critical speed, in revolutions per minute, for a ball mill 1600 mm in diameter charged with 100 mm balls......


a) 149


b) 276


c) 321


d) 419


c) NONE


1
Expert's answer
2021-08-25T15:16:01-0400

Critical speed for a ball mill

"N_c = \dfrac{1}{2π} (\dfrac{\sqrt{2g}}{(D-d)})"


"N_c = \dfrac1{2π}(\dfrac{\sqrt{2× 9.81}}{{1.6-0.1}})= 4.64\text{ rps}= 278.3\text{ rpm}"


b